Suppose that the perimeter of a rectangle is 24 inches and the length is twice the width. What are the dimensions of this rectangle and what is the area?
Answer by mananth(16946)

width =x
length = 2x
2x+2*2x=24
4x+2x=24
6x=24
x=4 inches
length = 8 inches
Area = L *W
8*4=32 square inches